# Generative AI in Transportation Planning: A Survey

### v1

@@ Abstract
The integration of generative artificial intelligence (GenAI) into transportation planning has the potential to revolutionize tasks such as demand forecasting, infrastructure design, policy evaluation, and traffic simulation. However, there is a critical need for a systematic framework to guide the adoption of GenAI in this interdisciplinary domain. In this survey, we—a multidisciplinary team of researchers spanning computer science and transportation engineering—present the first comprehensive framework for leveraging GenAI in transportation planning. Specifically, we introduce a new taxonomy that categorizes existing applications and methodologies into two perspectives: transportation planning tasks and computational techniques. From the transportation planning perspective, we examine the role of GenAI in automating descriptive, predictive, generative, simulation, and explainable tasks to enhance mobility systems. From the computational perspective, we detail advancements in data preparation, domain-specific fine-tuning, and inference strategies such as retrieval-augmented generation and zero-shot learning tailored to transportation applications. Additionally, we address critical challenges, including data scarcity, explainability, bias mitigation, and the development of domain-specific evaluation frameworks that align with transportation goals like sustainability, equity, and system efficiency. This survey aims to bridge the gap between traditional transportation planning methodologies and modern AI techniques, fostering collaboration and innovation. By addressing these challenges and opportunities, we seek to inspire future research that ensures ethical, equitable, and impactful use of generative AI in transportation planning.
@@ 1 Introduction
Recent advancements in generative artificial intelligence (GenAI) have showcased transformative potential across a wide array of fields, including healthcare [1, 2], finance [3, 4], scientific discovery [5, 6], transportation [7, 8, 9], and education [10]. Transportation planning, in particular, stands to benefit greatly from these advancements, as GenAI—particularly large language models (LLMs)—offers tools to address complex challenges such as scenario generation [11], multimodal system optimization [12], stakeholder engagement [13] and assistance with data analysis. By synthesizing insights from heterogeneous and dynamic data sources, LLMs have become essential tools for advancing transportation research and practice [14].
Transportation planning is a systematic process of developing strategies to manage and enhance the movement of people and goods across various transportation systems while addressing long-term societal goals. This process integrates data-driven methodologies to balance efficiency, equity, and sustainability in mobility systems, accounting for multimodal networks, infrastructure needs, and policy constraints [15, 16]. Transportation planning encompasses activities such as demand forecasting [17], infrastructure design [18], traffic management[19], and public engagement [20]. Traditionally, transportation planning relied on expert-driven frameworks, where decision-makers analyzed travel patterns, forecasted demands, and designed solutions based on statistical models and simulation techniques [21]. However, these methods often struggled to manage the increasing scale and complexity of modern transportation systems, particularly in incorporating diverse data sources, addressing real-time dynamics, and generating adaptive solutions. They also struggle with the costs of technology-forward solutions with public agency budgets and the ability to pay salaries and attract and keep technically strong talent.
@@ 2.2 Evolution of AI in Transportation
The introduction of AI-assisted methodologies marked a significant leap forward. Machine learning and predictive models enabled more efficient analysis of structured datasets, such as traffic counts, weather data, and household travel diary surveys and transit onboard surveys. AI-assisted systems provided predictions and optimization suggestions for tasks like congestion forecasting [66], signal optimization [57, 67], and traffic rerouting [68]. However, these systems required substantial human intervention for retraining, parameter tuning, and validation, making them adaptive but still reliant on expert oversight.
@@ 2.3 Generative AI Models for Transportation
GANs and VAEs. GANs and VAEs are widely used for generating synthetic transportation data, such as traffic flows or multimodal network scenarios. GANs are particularly effective in creating high-quality, realistic datasets by training two networks—the generator and discriminator—against each other [74, 75]. For instance, GANs can simulate traffic patterns under rare conditions, such as extreme weather events, enabling planners to evaluate system resilience [76]. VAEs, on the other hand, excel in encoding transportation data into latent variables, providing interpretable representations that support efficient scenario generation and simulation [77].
@@
Recent studies have also emphasized the role of large language models (LLMs) in supporting generative AI tasks. LLMs enhance scenario generation, explainability, and interactive planning through advanced text generation and reasoning capabilities [91]. For instance, LLMs can generate detailed descriptions of infrastructure projects, analyze public sentiment on transportation policies, and simulate stakeholder discussions, bridging the gap between technical outputs and human decision-making processes.
@@ 3.1 Core Transportation Applications
Demand Forecasting applies generative AI to predict travel demand patterns across multimodal systems. Models can fine-tune origin-destination (OD) matrices to align with observed traffic counts [109, 7], estimate the adoption of shared mobility services [110, 111, 112], and simulate long-term demand shifts due to demographic or economic changes [70, 52, 113, 88, 114]. These forecasts provide insights into system bottlenecks, enabling planners to balance travel loads across transportation networks effectively.
@@ 3.2 Generative AI Methodologies
Figure 4: Generative AI techniques.
@@
Benchmark Datasets form the foundation for evaluating generative AI performance in tasks like OD calibration, traffic simulation, and infrastructure planning. Examples include LargeST [124] for traffic analysis and traffic assignment datasets [125]. These datasets provide high-resolution inputs and ground-truth benchmarks to validate AI-driven insights. However, building comprehensive knowledge bases for these models remains a significant challenge, particularly for forecasting, where uncertainty and dynamic variables like population growth, land-use changes, and evolving travel behaviors play key roles. Generative AI requires diverse, harmonized data inputs, such as historical traffic data, multimodal interactions, and prior model outcomes, but integrating these into a machine-readable format is time-intensive and complex. The lack of standardization, interoperability, and real-time updates further complicates the process, creating a bottleneck in AI adoption. Future advancements must focus on automated dataset synthesis, standardized modeling frameworks, and adaptive real-time updates to ensure that generative AI can address the intricate and evolving demands of transportation systems effectively.
Data Preprocessing Strategies ensure clean, balanced, and representative inputs for AI models. Transportation datasets often require manual or automated annotation of traffic patterns [126], infrastructure attributes [127], or user feedback [128]. Generative AI methods, such as LLM-based automated labeling [129, 130], address annotation challenges while preserving data quality. Data augmentation techniques, such as synthetic OD matrix generation [131, 132], further expand datasets to simulate rare or extreme transportation scenarios.
Fine-Tuning Techniques enable generative AI to adapt to transportation-specific tasks efficiently. Methods like Low-Rank Adaptation (LoRA) [133] optimize model performance while minimizing computational demands. For complex tasks like OD calibration, Chain-of-Thought prompting [115] enhances the model’s reasoning ability to align OD matrices with observed traffic counts.
@@
Advanced Inference Strategies improve generative AI’s adaptability to real-world transportation challenges. Retrieval-Augmented Generation (RAG) dynamically integrates external data sources, such as live sensor feeds or policy records, to produce accurate and context-aware predictions [136, 45, 137, 138]. By modeling the interplay between agents and their environment, Graph-RAG [139] has proven effective in multi-modal tasks like route planning [140], which is possible be generalize to congestion management and safety-critical decision-making. Besides, Self-Consistency Decoding reduces variability in generated outputs, stabilizing recommendations for tasks like traffic signal optimization or route planning [141].
@@ 4.1 Descriptive Tasks for Data Fusion and Analytics
  • Data Heterogeneity: Integrating data from varied formats, resolutions, and sources remains a technical hurdle [156]. Transportation data often comes from diverse sources like GPS devices, traffic sensors, user apps, and public databases, etc [157]. And each of them has its unique structure and metadata. Standardizing these data streams into a unified format while preserving their integrity and context is essential but complex.
@@ 4.2 Predictive Tasks in Transportation Planning
  • Explainability and Trust: The lack of explainability in generative AI models, especially on deep learning models [178], is a significant hurdle for their adoption in high-stakes decision-making processes within transportation planning. This opacity makes it difficult for planners to understand and trust the reasoning behind AI-generated forecasts.
@@ 4.4 Simulation Tasks for Traffic Dynamics and Mixed-Autonomy Systems
Generative AI-driven simulations are particularly valuable for studying mixed-autonomy systems, where human-driven and autonomous vehicles (AVs) coexist. For example, frameworks like CoMAL [115] use LLM agents to simulate collaborative strategies for traffic merging, lane switching, or intersection coordination. These simulations optimize system-level performance by dynamically assigning roles (e.g., leader, follower) to AVs, reducing congestion and stop-and-go waves.
@@
  • Computational cost: Another critical challenge is the computational cost associated with running high-resolution, real-time simulations at scale. Generative AI models integrated into digital twins or predictive traffic systems require significant computational resources to process diverse data streams, such as real-time sensor feeds, GPS traces, and dynamic OD matrices [199]. While digital twins offer an opportunity to mirror real-world transportation systems for testing interventions, their reliance on generative AI magnifies the resource demands, particularly for urban-scale or multimodal networks. Developing lightweight and computationally efficient models is crucial to addressing these scalability constraints.
@@ 4.5 Trustworthiness in Generative AI-Based Transportation
Security. The robustness of Diffusion Models and Large Language Models against malicious manipulation is crucial for their safe deployment, especially in transportation, where security breaches can lead to life-threatening consequences. Two prevalent attack vectors are adversarial attacks [205] and backdoor attacks [206]. Adversarial attacks exploit vulnerabilities by introducing imperceptible perturbations to input data, deceiving the model into making incorrect predictions. In autonomous driving systems, such perturbations could cause a vehicle to misinterpret traffic signs or road conditions, leading to accidents. Backdoor attacks, on the other hand, embed concealed triggers within the model, allowing adversaries to control its behavior upon activation. For example, malicious actors could exploit these vulnerabilities to manipulate traffic signal control systems, resulting in traffic congestion or collisions. These security threats can compromise model integrity, leading to unpredictable and potentially harmful consequences in transportation networks.
@@ 5.3 Broad Source of Dataset Collection
This subsection explores the fine-tuning of generative AI models for transportation planning applications, using Origin-Destination (OD) Demand Calibration as an example. OD calibration refers to the process of estimating travel demand between different geographic locations by adjusting OD matrices to match observed traffic counts. This task is fundamental in transportation systems, as accurate demand estimation directly influences infrastructure planning, traffic control strategies, and network optimization. OD matrices are critical inputs for modeling transportation demand and evaluating policy interventions, as they help planners understand mobility patterns and design efficient transportation systems. Specifically, we employ the Open Traffic Intelligence (OpenTI) dataset [7], which provides detailed traffic data for OD calibration tasks. OD calibration involves adjusting OD matrices to align with observed traffic counts, a critical problem in transportation systems where accurate demand estimation directly impacts infrastructure planning, traffic control, and network optimization.
Domain-Specific Dataset The OpenTI dataset [7] is an open-source resource specifically designed to support traffic analysis and calibration tasks. It includes traffic flow observations, OD matrices, road network structures, and simulated travel demands. OpenTI’s detailed data structure provides both: (1) Input: Raw OD matrices and observed traffic counts across multiple time intervals. (2) Output: A series of calibrated OD matrices that better align with real-world traffic conditions. OpenTI captures diverse scenarios, including traffic demand variations during peak and off-peak hours, enabling fine-tuned generative AI models to improve the accuracy and adaptability of OD demand estimation.
@@ 5.5 Inference with LLMs: Zero-Shot In-Context Learning
Zero-Shot Learning (ZSL) in Urban Studies. ZSL leverages the pretrained capabilities of LLMs to solve downstream tasks with only textual prompts and no additional training [270]. In city studies, such as transportation accessibility, ZSL enables the extraction of meaningful insights from LLMs by asking well-crafted prompts, even when ground-truth data is unavailable. This makes ZSL an efficient and scalable solution for analyzing cities that lack comprehensive datasets.
@@
Comparison with Traditional Methods. Traditional approaches to estimating public transportation accessibility rely on geographic information systems (GIS) and extensive field surveys, which are resource-intensive [273]. ZSL, in contrast, extracts latent knowledge from LLMs using contextual prompts, offering a lightweight alternative for cities with limited data. As demonstrated in Li et al. (2024), LLM-derived features often show strong correlations with real-world transportation metrics, highlighting the potential of ZSL as a viable estimation method.
@@ 5.6 Inference with Generative AI: Few-Shot In-Context Learning
Designing Effective Few-shot Examples. The effectiveness of few-shot learning heavily relies on the selection of representative examples. For the task of Multi-Agent Traffic Optimization, CoMAL leverages LLM-based multi-agent systems to address the coordination of connected autonomous vehicles (CAVs) in mixed-autonomy traffic, where both autonomous and human-driven vehicles coexist. Few-shot examples are embedded in prompts to enable the LLM to infer strategies for optimizing traffic flow without extensive retraining. In the CoMAL framework, prompts include task descriptions, environment perception, and a few relevant examples of agent roles, driving instructions, and reasoning steps. By analyzing these few-shot examples, the LLM coordinates actions between agents, mitigates stop-and-go traffic waves, and enhances overall system efficiency. Few-Shot Coordination in Traffic Merging Scenarios. In the Merge Scenario of CoMAL, vehicles entering from an on-ramp create disturbances in traffic flow on the main highway. Few-shot learning prompts are used to guide the LLM in role allocation and motion planning to reduce congestion and ensure smooth merging behavior. Below is an example of a structured prompt that integrates the traffic task, agent roles, and driving plans.

### v2

@@ Abstract
The integration of generative artificial intelligence (GenAI) into transportation planning has the potential to revolutionize tasks such as demand forecasting, infrastructure design, policy evaluation, and traffic simulation. However, there is a critical need for a systematic framework to guide the adoption of GenAI in this interdisciplinary domain. In this survey, we, a multidisciplinary team of researchers spanning computer science and transportation engineering, present the first comprehensive framework for leveraging GenAI in transportation planning. Specifically, we introduce a new taxonomy that categorizes existing applications and methodologies into two perspectives: transportation planning tasks and computational techniques. From the transportation planning perspective, we examine the role of GenAI in automating descriptive, predictive, generative simulation, and explainable tasks to enhance mobility systems. From the computational perspective, we detail advancements in data preparation, domain-specific fine-tuning, and inference strategies such as retrieval-augmented generation and zero-shot learning tailored to transportation applications. Additionally, we address critical challenges, including data scarcity, explainability, bias mitigation, and the development of domain-specific evaluation frameworks that align with transportation goals like sustainability, equity, and system efficiency. This survey aims to bridge the gap between traditional transportation planning methodologies and modern AI techniques, fostering collaboration and innovation. By addressing these challenges and opportunities, we seek to inspire future research that ensures ethical, equitable, and impactful use of generative AI in transportation planning.
@@ 1 Introduction
Recent advancements in generative artificial intelligence (GenAI) have showcased transformative potential across a wide array of fields, including healthcare [1, 2], finance [3, 4], scientific discovery [5, 6], transportation [7, 8, 9], and education [10]. Transportation planning, in particular, stands to benefit greatly from these advancements, as GenAI, particularly large language models (LLMs), offers tools to address complex challenges such as scenario generation [11], multimodal system optimization [12], stakeholder engagement [13] and assistance with data analysis. By synthesizing insights from heterogeneous and dynamic data sources, LLMs have become essential tools for advancing transportation research and practice [14].
Transportation planning is a systematic process of developing strategies to manage and enhance the movement of people and goods across various transportation systems while addressing long-term societal goals. This process integrates data-driven methodologies to balance efficiency, equity, and sustainability in mobility systems, accounting for multimodal networks, infrastructure needs, and policy constraints [15, 16]. Transportation planning encompasses activities such as demand forecasting [17], infrastructure design [18], traffic management[19], and public engagement [20]. Traditionally, transportation planning relied on expert-driven frameworks, where decision-makers analyzed travel patterns, forecasted demands, and designed solutions based on statistical models and simulation techniques [21]. However, these methods often struggle to manage the increasing scale and complexity of modern transportation systems, particularly in incorporating diverse data sources, addressing real-time dynamics, and generating adaptive solutions. They also struggle with the costs of technology-forward solutions with public agency budgets and the ability to pay salaries to keep technically strong talent.
@@ 2.2 Evolution of AI in Transportation
The introduction of AI-assisted methodologies marked a significant leap forward. Machine learning, such as predictive models, enabled more efficient analysis of structured datasets, such as traffic counts, weather data, and household travel diary surveys and transit onboard surveys. AI-assisted systems provided predictions and optimization suggestions for tasks like congestion forecasting [66], signal optimization [57, 67], and traffic rerouting [68]. However, these systems required substantial human intervention for retraining, parameter tuning, and validation, making them adaptive but still reliant on expert oversight.
@@ 2.3 Generative AI Models for Transportation
GANs and VAEs. GANs and VAEs are widely used for generating synthetic transportation data, such as traffic flows or multi-modal traffic network scenarios. GANs are particularly effective in creating high-quality, realistic datasets by training two networks—the generator and discriminator—against each other [74, 75]. For instance, GANs can simulate traffic patterns under rare conditions, such as extreme weather events, enabling planners to evaluate system resilience [76]. VAEs, on the other hand, excel in encoding transportation data into latent variables, providing interpretable representations that support efficient scenario generation and simulation [77].
@@
Recent studies have also emphasized the role of LLMs in supporting generative AI tasks. LLMs enhance scenario generation, explainability, and interactive planning through advanced text generation and reasoning capabilities [91]. For instance, LLMs can generate detailed descriptions of infrastructure projects, analyze public sentiment on transportation policies, and simulate stakeholder discussions, bridging the gap between technical outputs and human decision-making processes.
@@ 3.1 Core Transportation Applications
Demand Forecasting applies generative AI to predict travel demand patterns across multimodal systems. Models can fine-tune origin-destination (O-D) matrices to align with observed traffic counts [109, 7], estimate the adoption of shared mobility services [110, 111, 112], and simulate long-term demand shifts due to demographic or economic changes [70, 52, 113, 88, 114]. These forecasts provide insights into system bottlenecks, enabling planners to balance travel loads across transportation networks effectively.
@@ 3.2 Generative AI Methodologies
Figure 4: Generative AI techniques. The taxonomy is spread with three main categories, Models, Learning types, and Techniques, and each of the categories is followed with distinct aspects for comprehensive research domain coverage.
@@
Benchmark Datasets form the foundation for evaluating generative AI performance in tasks like O-D calibration, traffic simulation, and infrastructure planning. Examples include LargeST [124] for traffic analysis and traffic assignment datasets [125]. These datasets provide high-resolution inputs and ground-truth benchmarks to validate AI-driven insights. However, building comprehensive knowledge bases for these models remains a significant challenge. Particularly for forecasting tasks, where uncertainty and dynamic variables play key roles, such as: population growth, land-use changes, and evolving travel behaviors. Generative AI requires diverse, harmonized data inputs, such as historical traffic data, multimodal interactions, and prior model outcomes, but integrating these into a machine-readable format is time-intensive and complex. The lack of standardization, interoperability, and real-time updates further complicates the process, creating a bottleneck in AI adoption. Future advancements must focus on automated dataset synthesis, standardized modeling frameworks, and adaptive real-time updates to ensure that generative AI can address the intricate and evolving demands of transportation systems effectively.
Data Preprocessing Strategies ensure clean, balanced, and representative inputs for AI models. Transportation datasets often require manual or automated annotation of traffic patterns [126], infrastructure attributes [127], or user feedback [128]. Generative AI methods, such as LLM-based automated labeling [129, 130], address annotation challenges while preserving data quality. Data augmentation techniques, such as synthetic O-D matrix generation [131, 132], further expand datasets to simulate rare or extreme transportation scenarios.
Fine-Tuning Techniques enable generative AI to adapt to transportation-specific tasks efficiently. Methods like Low-Rank Adaptation (LoRA) [133] optimize model performance while minimizing computational demands. For complex tasks like O-D calibration, Chain-of-Thought prompting [115] enhances the model’s reasoning ability to align O-D matrices with observed traffic counts.
@@
Advanced Inference Strategies improve generative AI’s adaptability to real-world transportation challenges. Retrieval-Augmented Generation (RAG) dynamically integrates external data sources, such as live sensor feeds or policy records, to produce accurate and context-aware predictions [136, 45, 137, 138]. By modeling the interplay between agents and their environment, Graph-RAG [139] has proven effective in multi-modal tasks like route planning [140], which can possibly be generalized to congestion management and safety-critical decision-making. Besides, Self-Consistency Decoding reduces variability in generated outputs, stabilizing recommendations for tasks like traffic signal optimization or route planning [141].
@@ 4.1 Descriptive Tasks for Data Fusion and Analytics
  • Data Heterogeneity: Integrating data from varied formats, resolutions, and sources remains a technical hurdle [156]. Transportation data often comes from diverse sources like GPS devices, traffic sensors, user apps, and public databases [157]. And each of them has its unique structure and metadata. Standardizing these data streams into a unified format while preserving their integrity and context is essential but complex.
@@ 4.2 Predictive Tasks in Transportation Planning
  • Explainability and Trust: The lack of explainability in generative AI models, especially in deep learning models [178], is a significant hurdle for their adoption in high-stakes decision-making processes within transportation planning. This opacity makes it difficult for planners to understand and trust the reasoning behind AI-generated forecasts.
@@ 4.4 Simulation Tasks for Traffic Dynamics and Mixed-Autonomy Systems
Generative AI-driven simulations are particularly valuable for studying mixed-autonomy systems, where human-driven vehicles (HVs) and autonomous vehicles (AVs) coexist. For example, frameworks like CoMAL [115] use LLM agents to simulate collaborative strategies for traffic merging, lane switching, or intersection coordination. These simulations optimize system-level performance by dynamically assigning roles (e.g., leader, follower) to AVs, reducing congestion and stop-and-go waves.
@@
  • Computational cost: Another critical challenge is the computational cost associated with running high-resolution, real-time simulations at scale. Generative AI models integrated into digital twins or predictive traffic systems require significant computational resources to process diverse data streams, such as real-time sensor feeds, GPS traces, and dynamic O-D matrices [199]. While digital twins offer an opportunity to mirror real-world transportation systems for testing interventions, their reliance on generative AI magnifies the resource demands, particularly for urban-scale or multimodal networks. Developing lightweight and computationally efficient models is crucial to for scalability constraints.
@@ 4.5 Trustworthiness in Generative AI-Based Transportation
Security. The robustness of Diffusion Models and LLM against malicious manipulation is crucial for their safe deployment, especially in transportation, where security breaches can lead to life-threatening consequences. Two prevalent attack vectors are adversarial attacks [205] and backdoor attacks [206]. Adversarial attacks exploit vulnerabilities by introducing imperceptible perturbations to input data, deceiving the model into making incorrect predictions. In autonomous driving systems, such perturbations could cause a vehicle to misinterpret traffic signs or road conditions, leading to accidents. Backdoor attacks, on the other hand, embed concealed triggers within the model, allowing adversaries to control its behavior upon activation. For example, malicious actors could exploit these vulnerabilities to manipulate traffic signal control systems, resulting in traffic congestion or collisions. These security threats can compromise model integrity, leading to unpredictable and potentially harmful consequences in transportation networks.
@@ 5.3 Broad Source of Dataset Collection
This subsection explores the fine-tuning of generative AI models for transportation planning applications, using Origin-Destination (O-D) Demand Calibration as an example. O-D calibration refers to the process of estimating travel demand between different geographic locations by adjusting O-D matrices to match observed traffic counts. This task is fundamental in transportation systems, as accurate demand estimation directly influences infrastructure planning, traffic control strategies, and network optimization. O-D matrices are critical inputs for modeling transportation demand and evaluating policy interventions, as they help planners understand mobility patterns and design efficient transportation systems. Specifically, we employ the Open Traffic Intelligence (OpenTI) dataset [7], which provides detailed traffic data for O-D calibration tasks. O-D calibration involves adjusting O-D matrices to align with observed traffic counts, a critical problem in transportation systems where accurate demand estimation directly impacts infrastructure planning, traffic control, and network optimization.
Domain-Specific Dataset The OpenTI dataset [7] is an open-source resource specifically designed to support traffic analysis and calibration tasks. It includes traffic flow observations, O-D matrices, road network structures, and simulated travel demands. OpenTI’s detailed data structure provides both: (1) Input: Raw OD matrices and observed traffic counts across multiple time intervals. (2) Output: A series of calibrated OD matrices that better align with real-world traffic conditions. OpenTI captures diverse scenarios, including traffic demand variations during peak and off-peak hours, enabling fine-tuned generative AI models to improve the accuracy and adaptability of OD demand estimation.
@@ 5.5 Inference with LLMs: Zero-Shot In-Context Learning
Zero-Shot Learning (ZSL) in Urban Studies. ZSL leverages the pre-trained capabilities of LLMs to solve downstream tasks with only textual prompts and no additional training [270]. In city studies, such as transportation accessibility, ZSL enables the extraction of meaningful insights from LLMs by asking well-crafted prompts, even when ground-truth data is unavailable. This makes ZSL an efficient and scalable solution for analyzing cities that lack comprehensive datasets.
@@
Comparison with Traditional Methods. Traditional approaches to estimating public transportation accessibility rely on geographic information systems (GIS) and extensive field surveys, which are resource-intensive [273]. ZSL, in contrast, extracts latent knowledge from LLMs using contextual prompts, offering a lightweight alternative for cities with limited data. As demonstrated in [jin2023large, 116, li2024can], LLM-derived features often show strong correlations with real-world transportation metrics, highlighting the potential of ZSL as a viable estimation method.
@@ 5.6 Inference with Generative AI: Few-Shot In-Context Learning
Designing Effective Few-shot Examples. The effectiveness of few-shot learning heavily relies on the selection of representative examples. For the task of Multi-Agent Traffic Optimization, CoMAL leverages LLM-based multi-agent systems to address the coordination of connected autonomous vehicles (CAVs) in mixed-autonomy traffic, where both autonomous and human-driven vehicles coexist. Few-shot examples are embedded in prompts to enable the LLM to infer strategies for optimizing traffic flow without extensive retraining. In the CoMAL framework, prompts include task descriptions, environment perception, and a few relevant examples of agent roles, driving instructions, and reasoning steps. By analyzing these few-shot examples, the LLM ***For details of LLM version, please visit original paper [115]. coordinates actions between agents, mitigates stop-and-go traffic waves, and enhances overall system efficiency. Few-Shot Coordination in Traffic Merging Scenarios. In the Merge Scenario of CoMAL, vehicles entering from an on-ramp create disturbances in traffic flow on the main highway. Few-shot learning prompts are used to guide the LLM in role allocation and motion planning to reduce congestion and ensure smooth merging behavior. Below is an example of a structured prompt that integrates the traffic task, agent roles, and driving plans.